# Information Density Bounds for Privacy

### v1

@@ V-B Secrets with General Alphabets
Now, we consider the more general case where $(X,Y)$ takes values in the product space $(\mathcal{X}\times\mathcal{Y},\mathcal{S}_{\mathcal{X}}\otimes\mathcal{S}_{\mathcal{Y}})$. Note that the first statement in Theorem 4 has no useful parallel for secrets with infinite alphabets due to its dependence on $p_{\mathrm{min}}$. Nonetheless, below we provide a more general version of the second statement in Theorem 4. Its proof is provided in Appendix G.
@@
Suppose $P_{XY}\ll P_{X}\times P_{Y}$. If $\Lambda(X\to Y)$ is essentially bounded, then $\ell(X\to Y)$ is also essentially bounded.
Note that the opposite direction in the above result does not hold. For example, while PML is always bounded for finite random variables, PMC can be infinite even in these cases. Hence, Theorem 5 demonstrates that PMC is a stricter measure compared to PML. Interestingly, below we show that an even stronger statement is possible.
Theorem 6.
$\Lambda(X\to Y)$ is essentially bounded if and only if the privacy mechanism $P_{Y\mid X}$ satisfies $\varepsilon$-LDP for some $0\leq\varepsilon<\infty$.
The above result is proved in Appendix H. Note that Theorem 5 may be viewed as a consequence of Theorem 6 since a mechanism satisfying LDP has bounded PML.
@@ VI Discussion: PMC As an Information Leakage Measure
[…]
When $X$ is a finite random variable, the initial uncertainty in $X$ is finite according to all the standard measures of uncertainty such as the Rényi entropies. Therefore, a direct consequence of (48) is that information leaked about finite random variables should invariably remain finite, regardless of the mechanism used. This finiteness criterion may be viewed as an axiomatic requirement for information leakage measures and is satisfied by most measures such as PML, mutual information, and $f$-information, among others.
@@
A second and more practical limitation of PMC is that while it is a context-aware privacy measure, its application is effectively restricted to the class of mechanisms compatible with LDP and DP. In finite settings, this is evident from Theorem 4 where PMC bounds can be derived only in the high-privacy regime. This trend also extends to the general case as indicated by Theorem 6. These results suggest that PMC has a more limited application compared to PML and is more restrictive in terms of utility.
As a concrete example, consider the task of estimating the mean of a Gaussian random variable by adding Gaussian noise to the sample mean. Such a problem is well-defined under PML since PML allows the addition of Gaussian noise to a Gaussian random variable (see [13, Example 4]). However, this problem cannot be studied with PMC. As shown in Example 6 and specifically (40), the lower bound on PMC becomes infinite when Gaussian noise is added to an unbounded random variable. This observation aligns with the sensitivity-based analysis in DP, where the sample mean’s sensitivity becomes infinite when the data domain is unbounded.
That being said, while using DP and LDP mechanisms, allowing higher values of PMC compared to PML can result in improved utility. In this context, employing a framework such as ALIP offers better flexibility and more favorable privacy-utility tradeoffs compared to LIP.
@@ VII Conclusions
This work further explored the relationship between the upper and lower bounds on the information density by examining how PMC and PML guarantees relate. Our findings reveal that PMC is a more stringent privacy measure and its application is limited to mechanisms satisfying differential privacy-type guarantees. Notably, PMC can be infinite even when applied to finite random variables. This characteristic, which sets PMC apart from other measures such as PML and $f$-information, arises from the definition of PMC: If the posterior expected value of a cost function is zero then PMC is infinite.
Overall, our work significantly bridges the gaps in understanding the relationships between various privacy frameworks. Looking ahead, the findings in this paper offer crucial insights for selecting the appropriate framework to ensure privacy in different applications. If the objective is to adopt a conservative approach that prevents any definitive inference about the secret, then LDP, LIP, and ALIP would be suitable choices. When information about the prior distribution is available, LIP and ALIP are preferred as they enable higher utility compared to LDP. ALIP provides additional flexibility by allowing distinct upper and lower bounds, while LIP simplifies the design by imposing symmetric bounds. Conversely, when the prior is not known, then LDP would be a suitable choice since it offers equally strong protection regardless of the prior. In addition, PML emerges as a particularly fitting choice when working with mechanisms that do not satisfy privacy in the traditional differential privacy sense, for instance, mechanisms with unbounded domains. PML provides the flexibility to tune the privacy parameter to achieve different types of protection, from preventing any definitive inference about the secret to high-utility mechanisms that hide only the least likely and atypical features of the secret [21].
Finally, our work has practical implications for designing privacy mechanisms since a mechanism initially designed to meet a specific privacy guarantee may also satisfy other types of guarantees. Using the relations between various measures presented here, it is possible to examine if a mechanism provides other guarantees beyond its original design.
@@ Appendix A Proof of Theorem 2
Fix $y\in\mathcal{Y}$. First, we argue that given a $U$ satisfying the Markov chain $U-X-Y$ there exists a non-negative cost function $c_{U}$ such that $\Lambda_{U}(X\to y)=\Lambda_{c}(X\to y)$. Given $U$, let the cost function $c_{U}:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{+}$ be defined as $c_{U}(x,u)=1-P_{U\mid X=x}(u)$ with $x\in\mathcal{X}$ and $u\in\mathcal{U}$. Then, we have
@@
Now, we show that for each non-negative cost function $c$, there exists $U$ satisfying the Markov chain $U-X-Y$ such that $\Lambda_{c}(X\to y)=\Lambda_{U}(X\to y)$. Fix a cost function $c$. Without loss of generality, we assume that $0\leq c(x,w)\leq 1$ for all $x\in\mathcal{X}$ and $w\in\mathcal{W}$. This can be achieved through normalizing $c$ by $\max_{x,w}c(x,w)$.
@@
$=\min\Bigg{\{}\sum_{x}\Big{(}1-\delta\frac{c(x,w_{S})}{k}-(1-\delta)\frac{c(x,w_{T})}{k}\Big{)}P_{X}(x),\sum_{x}\Big{(}\delta c(x,w_{S})+(1-\delta)c(x,w_{T})\Big{)}P_{X}(x)\Bigg{\}}$
@@
$\inf_{P_{\hat{U}\mid Y}}\;\mathbb{P}[U_{\delta}\neq\hat{U}\mid Y=y]=\min_{u\in[k+1]}\sum_{x}(1-P_{U_{\delta}\mid X=x}(u))P_{X\mid Y=y}(x)$
$=\min\Bigg{\{}\sum_{x}\Big{(}1-\delta\frac{c(x,w_{S})}{k}-(1-\delta)\frac{c(x,w_{T})}{k}\Big{)}P_{X\mid Y=y}(x),\sum_{x}\Big{(}\delta c(x,w_{S})+(1-\delta)c(x,w_{T})\Big{)}P_{X\mid Y=y}(x)\Bigg{\}}$
@@
$\exp\Big{(}\Lambda_{U_{\delta}}(X\to y)\Big{)}=\frac{\sum_{x}\Big{(}\delta c(x,w_{S})+(1-\delta)c(x,w_{T})\Big{)}P_{X}(x)}{\sum_{x}\Big{(}\delta c(x,w_{S})+(1-\delta)c(x,w_{T})\Big{)}P_{X\mid Y=y}(x)}.$
@@
==$\infty$,= ∞ ,
since $10<P_{X}(\mathcal{E})<$ 1.
@@ Appendix B Proof of Theorem 3
Fix $y\in\mathcal{Y}$. The proof follows along the same lines of the proof of [13, Thm. 3]. We begin by noting that since $(\mathcal{X},\mathcal{S}_{\mathcal{X}})$ is standard Borel, then $P_{XY}$ can be disintegrated into the marginal $P_{Y}$ and a conditional distribution $P_{X\mid Y}$ [59, Thm. IV.2.18]. We use this fact to simplify the denominator of (26). The following lemma is proved in Appendix C.
@@
Of course, $\mathcal{D}_{w}^{\varepsilon}\in\mathcal{S}_{\mathcal{X}}$ for all $w\in\mathcal{W}$. Fix a large constant $M>0$ and consider the cost function $c^{*}:\mathcal{X}\times\mathcal{W}\to\mathbb{R}_{+}$ defined by
[…]
@@
$=\frac{\inf\limits_{w\in\mathcal{W}:P_{X}(\mathcal{D}_{w}^{\varepsilon})>0}\frac{P_{X}(\mathcal{D}_{w}^{\varepsilon})}{P_{X}(\mathcal{D}_{w}^{\varepsilon})}}{\inf\limits_{w\in\mathcal{W}:P_{X}(\mathcal{D}_{w}^{\varepsilon})>0}\frac{P_{X\mid Y=y}(\mathcal{D}_{w}^{\varepsilon})}{P_{X}(\mathcal{D}_{w}^{\varepsilon})}}$ (52b)
@@
where (52b) follows by taking $M$ to be sufficiently large and
[…]
@@ Appendix C Proof of Lemma 2
$\delta_{w}(\mathcal{A})=\begin{cases}1&\mathrm{if}\;w\in\mathcal{A},\\ 0&\mathrm{if}\;w\notin\mathcal{A},\\ \end{cases}$
@@
$\inf_{P_{W\mid Y}}\mathbb{E}\left[c(X,W)\mid Y=y\right]$ $\leq\int_{\mathcal{X}}P_{X\mid Y=y}(dx)\int_{\mathcal{W}}c(x,w)\;\delta_{w^{\prime}}(dw)$
$=\int_{\mathcal{X}}c(x,w^{\prime})\;P_{X\mid Y=y}(dx)$
$\leq a.$
@@ Appendix D Proof of Lemma 1
    $\Lambda(X_{1},\ldots,X_{k}\to y_{1},\ldots,y_{k})$ $=D_{\infty}(P_{X_{1},\ldots,X_{k}}\|P_{X_{1},\ldots,X_{k}\mid Y_{1}=y_{1},\ldots Y_{k}=y_{k}})$ (54)
@@
    $=\log\;\frac{\theta\int_{\mathcal{X}}f_{Y\mid X=x}(y)f_{X}(x)\;\mu(dx)+(1-\theta)\int_{\mathcal{X}}f_{Y\mid X=x}(y)g_{X}(x)\;\mu(dx)}{\operatorname*{ess\,inf}\limits_{P_{\theta}}\;f_{Y\mid X}(y\mid X)}$ (59)
    $=\log\;\frac{\theta f_{Y}(y)+(1-\theta)g_{Y}(y)}{\operatorname*{ess\,inf}\limits_{P_{\theta}}\;f_{Y\mid X}(y\mid X)}$ (60)
    $\geq\log\;\left(\theta\frac{f_{Y}(y)}{\operatorname*{ess\,inf}\limits_{P_{X}}\;f_{Y\mid X}(y\mid X)}+(1-\theta)\frac{g_{Y}(y)}{\operatorname*{ess\,inf}\limits_{Q_{X}}\;f_{Y\mid X}(y\mid X)}\right)$ (61)
    $\geq\theta\log\;\frac{f_{Y}(y)}{\operatorname*{ess\,inf}\limits_{P_{X}}\;f_{Y\mid X}(y\mid X)}+(1-\theta)\log\;\frac{g_{Y}(y)}{\operatorname*{ess\,inf}\limits_{Q_{X}}\;f_{Y\mid X}(y\mid X)}$ (62)
@@
  5. 5.
    Fix $y\in\mathcal{Y}$. Consider the Markov chain $Z-X-Y$. The kernel $P_{Z\mid X}$ induces distributions $P_{Z}=P_{Z\mid X}\circ P_{X}$ and $P_{Z\mid Y=y}=P_{Z\mid X}\circ P_{X\mid Y=y}$. Therefore, we can write
    $\Lambda(Z\to y)$ $=D_{\infty}(P_{Z}\|P_{Z\mid Y=y})=D_{\infty}(P_{Z\mid X}\circ P_{X}\|P_{Z\mid X}\circ P_{X\mid Y=y})$
@@
  6. 6.
    Consider the Markov chain $X-Y-Z$. The kernel $P_{Z\mid Y}$ induces distributions $P_{XZ}=P_{Z\mid Y}\circ P_{XY}$ and $P_{Z}=P_{Z\mid Y}\circ P_{Y}$. Hence, we can write
    $\operatorname*{ess\,sup}_{P_{Z}}\Lambda(X\to Z)$ $=D_{\infty}(P_{X}\times P_{Z}\|P_{XZ})=D_{\infty}(P_{X}\times(P_{Z\mid Y}\circ P_{Y})\|P_{Z\mid Y}\circ P_{XY})$
    $=D_{\infty}(P_{Z\mid Y}\circ(P_{X}\times P_{Y})\|P_{Z\mid Y}\circ P_{XY})\leq D_{\infty}(P_{X}\times P_{Y}\|P_{XY})$
@@
    $\leq\log\;\operatorname*{ess\,sup}_{P_{X}}\;\frac{f_{Y_{1}}(y_{1})}{f_{Y_{1}\mid X}(y_{1}\mid X)}+\log\;\operatorname*{ess\,sup}_{P_{X}}\;\frac{f_{Y_{2}\mid Y_{1}}(y_{2}\mid y_{1})}{f_{Y_{2}\mid Y_{1},X}(y_{2}\mid y_{1},X)}$ (67)
    $\leq\log\;\operatorname*{ess\,sup}_{P_{X}}\;\frac{f_{Y_{1}}(y_{1})}{f_{Y_{1}\mid X}(y_{1}\mid X)}+\log\;\operatorname*{ess\,sup}_{P_{X\mid Y_{1}=y_{1}}}\;\frac{f_{Y_{2}\mid Y_{1}}(y_{2}\mid y_{1})}{f_{Y_{2}\mid Y_{1},X}(y_{2}\mid y_{1},X)},$ (68)

### v2

@@ V-B Secrets with General Alphabets
Now, we consider the more general case where $(X,Y)$ takes values in the product space $(\mathcal{X}\times\mathcal{Y},\mathcal{S}_{\mathcal{X}}\otimes\mathcal{S}_{\mathcal{Y}})$. Earlier, we discussed that a mechanism can have finite PML but infinite PMC. This invites the question: Is PMC inherently more restrictive than PML? That is, does bounded PMC imply bounded PML? The following two examples show that, in general, the answer is no.
Example 8.
Let $X$ and $Y$ be countably infinite random variables with alphabets $\mathcal{X}=\mathcal{Y}=\{0,1,\ldots\}$. Consider the prior
$P_{X}(0)=\frac{1}{2},\quad P_{X}(n)=3^{-n}\text{ for }n\geq 1\\$
and the mechanism
$P_{Y\mid X=0}(m)=\begin{cases}\frac{3}{4}&\text{if }m=0,\\ \frac{3^{-m}}{2}&\text{if }m\geq 1,\end{cases}$
$P_{Y\mid X=n}(m)=\begin{cases}\frac{1}{4}&\text{if }m=0,n\geq 1\\ \frac{1}{2}+\frac{3^{-m}}{2}&\text{if }m,n\geq 1,m=n,\\ \frac{3^{-m}}{2}&\text{if }m,n\geq 1,m\neq n.\end{cases}$
Using the above values, it can be verified that $Y$ has the same marginal distribution as $X$, that is,
$P_{Y}(0)=\frac{1}{2},\quad P_{Y}(m)=3^{-m}\text{ for }m\geq 1.$
We now calculate the PML of the above mechanism at $Y=m\geq 1$. We have
$\ell(X\to m)$ $=\log\,\frac{\sup_{n}P_{Y\mid X=n}(m)}{P_{Y}(m)}$
$=\log\,\frac{\frac{1}{2}+\frac{3^{-m}}{2}}{3^{-m}}$
$=\log\!\Big(\frac{3^{m}+1}{2}\Big).$
Thus, $\ell(X\to m)\to\infty$ as $m\to\infty$ and the PML is unbounded.
Next, we calculate the PMC. For $m\geq 1$ we have
$\Lambda(X\to m)=\log\,\frac{P_{Y}(m)}{\inf_{n}P_{Y\mid X=n}(m)}=\log\,\frac{3^{-m}}{\frac{3^{-m}}{2}}=\log 2,$
and for $m=0$ we have
$\Lambda(X\to 0)=\log\,\frac{P_{Y}(0)}{\inf_{n}P_{Y\mid X=n}(0)}=\log 2.$
Therefore, the mechanism satisfies $\log 2$-PMC.
We also construct an example with absolutely continuous random variables.
Example 9.
Let
$c_{1}(x,y)=2(xy)^{-2}\Big(\frac{1}{x}+\frac{1}{y}-1\Big)^{-3},\quad x,y\in(0,1),$
denote the density of the Clayton copula $c_{\theta}(\cdot,\cdot)$ with parameter $\theta=1$ [11, 47]. A copula defines a joint distribution with uniform marginals. That is, $X\sim\mathsf{U}(0,1)$ and $Y\sim\mathsf{U}(0,1)$ since
$\int_{0}^{1}c_{1}(x,y)\,dx=1\;\text{ for all }y\in(0,1),$
$\int_{0}^{1}c_{1}(x,y)\,dy=1\;\text{ for all }x\in(0,1).$
Observe that $c_{1}(t,t)\to\infty$ when $t\to 0$, so the density is unbounded.
Now, fix $\eta\in(0,1)$ and consider the following joint density for $X$ and $Y$:
$f_{XY}(x,y)=\eta+(1-\eta)\,c_{1}(x,y),\quad x,y\in(0,1).$
This mixture preserves the marginal distributions, so $X\sim\mathsf{U}(0,1)$ and $Y\sim\mathsf{U}(0,1)$ still hold. Furthermore, the density remains unbounded above, but it is now bounded below by $\eta$.
We now compute both the PML and the PMC. The largest PML over the outcomes $y\in(0,1)$ is
$\sup_{y\in(0,1)}\ell(X\to y)$ $=\log\,\sup_{x,y\in(0,1)}\,\frac{f_{XY}(x,y)}{f_{X}(x)f_{Y}(y)}$
$\geq\log\,\sup_{t\in(0,1)}f_{XY}(t,t)=\infty.$
Hence, the PML is unbounded. On the other hand, for the PMC, we have
$\sup_{y\in(0,1)}\Lambda(X\to y)=\log\,\sup_{x,y\in(0,1)}\,\frac{f_{X}(x)f_{Y}(y)}{f_{XY}(x,y)}\leq\log\frac{1}{\eta},$
so the mechanism satisfies $\log\frac{1}{\eta}$-PMC.
Together, these two examples demonstrate that a mechanism can have bounded PMC but unbounded PML.
Lastly, we show that the connections between LDP, PML, and PMC established in Proposition 1, Proposition 3, and Corollary 3 partially extend to general alphabets. The proof is given in Appendix H.
@@
Let $P_{XY}=P_{Y\mid X}\times P_{X}$ be the joint distribution of $X$ and $Y$. If the mechanism $P_{Y\mid X}$ satisfies $\varepsilon$-LDP with $\varepsilon\geq 0$, then it satisfies $\varepsilon$-PMC and $\varepsilon$-PML. Conversely, if the mechanism satisfies $\varepsilon_{1}$-PMC and $\varepsilon_{2}$-PML, with $\varepsilon_{1},\varepsilon_{2}\geq 0$, then it satisfies $\varepsilon_{1}+\varepsilon_{2}$-LDP.
The above result demonstrates that, in general, the class of LDP mechanisms coincides with the class of mechanisms that have both bounded PMC and bounded PML. That is, LDP is equivalent to simultaneously restricting the threat imposed by risk-averse and opportunistic adversaries. This can be seen as a general-alphabet extension of [71, Prop. 1].
@@ VI Discussion: PMC As an Information Leakage Measure
$\text{information leakage}=\\ \text{initial uncertainty}-\text{remaining uncertainty}.$ (55)
When $X$ is a finite random variable, the initial uncertainty in $X$ is finite according to all the standard measures of uncertainty, such as the Rényi entropies. Therefore, a direct consequence of (55) is that information leaked about finite random variables should invariably remain finite, regardless of the mechanism used. This finiteness criterion may be viewed as an axiomatic requirement for information leakage measures and is satisfied by most measures, such as PML, mutual information, and $f$-information, among others.
@@
A second consequence of the different threat models is that the class of mechanisms satisfying PMC and PML guarantees, while overlapping, are distinct. When the underlying alphabets are finite, PMC is strictly stronger since a bound on PMC always implies a bound on PML, but the reverse need not hold. In general measurable spaces, the connection is more nuanced. A mechanism can have bounded PMC, bounded PML, both, or neither. Moreover, as long as $P_{XY}\ll P_{X}\times P_{Y}$, PML is well-defined even if it is unbounded. However, this is not the case for PMC, which may be undefined altogether. As a concrete example, consider the task of privately estimating the mean of a Gaussian random variable by adding Gaussian noise to the sample mean. This setup is well-defined under PML, since PML permits the addition of Gaussian noise to a Gaussian random variable (see [54, Example 4]). In contrast, PMC cannot be defined in this setting. As shown in Example 7 and specifically in (46), the lower bound on PMC becomes infinite when Gaussian noise is added to an unbounded random variable.
@@ VII Conclusions
This work further examined the connections between upper and lower bounds on the information density by analyzing how PMC and PML relate. When applied to finite random variables, PMC emerges as a more stringent privacy measure, applicable only to mechanisms that satisfy differential privacy-type guarantees. This characteristic, which sets PMC apart from other measures such as PML and $f$-information, arises from the definition of PMC: If the posterior expected value of a cost function is zero, then PMC is infinite. In general measurable spaces, we showed that the class of mechanisms satisfying LDP is exactly the set of those with bounded PMC and bounded PML. Put differently, LDP is equivalent to simultaneously restricting the threat imposed by risk-averse and opportunistic adversaries.
Overall, our work significantly bridges the gaps in understanding the relationships between various privacy frameworks. Looking ahead, the findings in this paper offer crucial insights for selecting the appropriate framework to ensure privacy in different applications. If the objective is to adopt a conservative approach that prevents any definitive inference about the secret, then LDP, LIP, and ALIP would be suitable choices. When information about the prior distribution is available, LIP and ALIP are preferred as they enable higher utility compared to LDP. ALIP provides additional flexibility by allowing distinct upper and lower bounds, while LIP simplifies the design by imposing symmetric bounds. Conversely, when the prior is not known, then LDP would be a suitable choice since it offers equally strong protection regardless of the prior.
Finally, we highlight how the results in this work can inform the practical design of privacy mechanisms. A mechanism originally designed to satisfy one privacy guarantee may also fulfill others. Using the relationships between privacy measures developed in this paper, one can assess whether a given mechanism provides additional guarantees beyond its original design. In addition, many of the canonical examples studied here, such as the generalized randomized response and Laplace mechanisms, form the basis for more sophisticated algorithms used in applications like federated learning, data sharing, and privacy-preserving statistics. By introducing a context-aware measure that captures a notion of privacy complementary to existing ones, this work takes a step toward designing high-utility systems with strong and meaningful guarantees.
@@ Appendix A Proof of Theorem 3
Fix $y\in\mathcal{Y}$. First, we argue that given a $U$ satisfying the Markov chain UXYU-X-Y there exists a non-negative cost function $c_{U}$ such that $\Lambda_{U}(X\to y)=\Lambda_{c}(X\to y)$. Given $U$, let the cost function $c_{U}:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{+}$ be defined as $c_{U}(x,u)=1-P_{U\mid X=x}(u)$ with $x\in\mathcal{X}$ and $u\in\mathcal{U}$. Then, we have
@@
Now, we show that for each non-negative cost function $c$, there exists $U$ satisfying the Markov chain UXYU-X-Y such that $\Lambda_{c}(X\to y)=\Lambda_{U}(X\to y)$. Fix a cost function $c$. Without loss of generality, we assume that $0\leq c(x,w)\leq 1$ for all $x\in\mathcal{X}$ and $w\in\mathcal{W}$. This can be achieved through normalizing $c$ by $\max_{x,w}c(x,w)$.
@@
$=\min\Bigg\{\sum_{x}\Big(1-\delta\frac{c(x,w_{S})}{k}-(1-\delta)\frac{c(x,w_{T})}{k}\Big)P_{X}(x),$
$\hskip 60.00009pt\sum_{x}\Big(\delta c(x,w_{S})+(1-\delta)c(x,w_{T})\Big)P_{X}(x)\Bigg\}$
@@
$\inf_{P_{\hat{U}\mid Y}}\;\mathbb{P}[U_{\delta}\neq\hat{U}\mid Y=y]$
$=\min_{u\in[k+1]}\sum_{x}(1-P_{U_{\delta}\mid X=x}(u))P_{X\mid Y=y}(x)$
$=\min\!\Bigg\{\!\sum_{x}\!\Big(\!1-\delta\frac{c(x,w_{S})}{k}-(1-\delta)\frac{c(x,w_{T})}{k}\Big)P_{X\mid Y=y}(x),$
$\hskip 50.00008pt\sum_{x}\Big(\delta c(x,w_{S})+(1-\delta)c(x,w_{T})\Big)P_{X\mid Y=y}(x)\Bigg\}$
@@
$\exp\Big(\Lambda_{U_{\delta}}(X\to y)\Big)\\ =\frac{\sum_{x}\Big(\delta c(x,w_{S})+(1-\delta)c(x,w_{T})\Big)P_{X}(x)}{\sum_{x}\Big(\delta c(x,w_{S})+(1-\delta)c(x,w_{T})\Big)P_{X\mid Y=y}(x)}.$
@@
$=\infty,$
since $0<P_{X}(\mathcal{E})<1$.
@@ Appendix B Proof of Theorem 4
Fix $y\in\mathcal{Y}$. The proof follows along the same lines of the proof of [54, Thm. 3]. We begin by noting that since $(\mathcal{X},\mathcal{S}_{\mathcal{X}})$ is standard Borel, then $P_{XY}$ can be disintegrated into the marginal $P_{Y}$ and a conditional distribution $P_{X\mid Y}$ [10, Thm. IV.2.18]. We use this fact to simplify the denominator of (31). The following lemma is proved in Appendix C.
@@
Of course, $\mathcal{D}_{w}^{\varepsilon}\in\mathcal{S}_{\mathcal{X}}$ for all $w\in\mathcal{W}$. Fix a constant $M>1$ and consider the cost function $c^{*}:\mathcal{X}\times\mathcal{W}\to\mathbb{R}_{+}$ defined by
$c^{*}(x,w)=\begin{cases}\frac{1}{P_{X}(\mathcal{D}_{w}^{\varepsilon})}\,\mathbf{1}_{\mathcal{D}_{w}^{\varepsilon}}(x)&\mathrm{if}\;P_{X}(\mathcal{D}_{w}^{\varepsilon})>0,\\ M&\mathrm{if}\;P_{X}(\mathcal{D}_{w}^{\varepsilon})=0.\end{cases}$
@@
$=\frac{1}{\min\left\{M,\inf\limits_{w\in\mathcal{W}:P_{X}(\mathcal{D}_{w}^{\varepsilon})>0}\frac{P_{X\mid Y=y}(\mathcal{D}_{w}^{\varepsilon})}{P_{X}(\mathcal{D}_{w}^{\varepsilon})}\right\}}$ (59b)
$\geq\frac{1}{\inf\limits_{w\in\mathcal{W}:P_{X}(\mathcal{D}_{w}^{\varepsilon})>0}\frac{P_{X\mid Y=y}(\mathcal{D}_{w}^{\varepsilon})}{P_{X}(\mathcal{D}_{w}^{\varepsilon})}}$
@@
where (59b) follows because $M>1$, and
$\bar{f}(x)\coloneqq\sum_{w\in\mathcal{W}}\frac{P_{X}(\mathcal{D}_{w}^{\varepsilon})}{P_{X\mid Y=y}(\mathcal{D}_{w}^{\varepsilon})}\;\mathbf{1}_{\mathcal{D}_{w}^{\varepsilon}}(x),\quad x\in\mathcal{X}.$
@@ Appendix C Proof of Lemma 2
[…]
@@
$\inf_{P_{W\mid Y}}\mathbb{E}$ $\left[c(X,W)\mid Y=y\right]$
$\leq\int_{\mathcal{X}}P_{X\mid Y=y}(dx)\int_{\mathcal{W}}c(x,w)\;\delta_{w^{\prime}}(dw)$
$=\int_{\mathcal{X}}c(x,w^{\prime})\;P_{X\mid Y=y}(dx)\leq a.$
@@ Appendix D Proof of Lemma 1
    $\Lambda(X_{1},\ldots,$ $X_{k}\to y_{1},\ldots,y_{k})$ (61)
    $=D_{\infty}(P_{X_{1},\ldots,X_{k}}\|P_{X_{1},\ldots,X_{k}\mid Y_{1}=y_{1},\ldots Y_{k}=y_{k}})$ (62)
@@
    $=\log\Bigg(\frac{\theta\int f_{Y\mid X=x}(y)f_{X}(x)\mu(dx)}{\operatorname*{ess\,inf}\limits_{P_{\theta}}\;f_{Y\mid X}(y\mid X)}$ (67)
    $\hskip 70.0001pt+\frac{(1-\theta)\int f_{Y\mid X=x}(y)g_{X}(x)\mu(dx)}{\operatorname*{ess\,inf}\limits_{P_{\theta}}\;f_{Y\mid X}(y\mid X)}\Bigg)$ (68)
    $=\log\;\Bigg(\frac{\theta f_{Y}(y)}{\operatorname*{ess\,inf}\limits_{P_{\theta}}\;f_{Y\mid X}(y\mid X)}+\frac{(1-\theta)g_{Y}(y)}{\operatorname*{ess\,inf}\limits_{P_{\theta}}\;f_{Y\mid X}(y\mid X)}\Bigg)$ (69)
    $\geq\log\Bigg(\theta\frac{f_{Y}(y)}{\operatorname*{ess\,inf}\limits_{P_{X}}f_{Y\mid X}(y\mid X)}$ (70)
    $\hskip 80.00012pt+(1-\theta)\frac{g_{Y}(y)}{\operatorname*{ess\,inf}\limits_{Q_{X}}f_{Y\mid X}(y\mid X)}\Bigg)$ (71)
    $\geq\theta\log\;\frac{f_{Y}(y)}{\operatorname*{ess\,inf}\limits_{P_{X}}\;f_{Y\mid X}(y\mid X)}$ (72)
    $\hskip 70.0001pt+(1-\theta)\log\;\frac{g_{Y}(y)}{\operatorname*{ess\,inf}\limits_{Q_{X}}\;f_{Y\mid X}(y\mid X)}$ (73)
@@
  5. (v)
    Fix $y\in\mathcal{Y}$. Consider the Markov chain ZXYZ-X-Y. The kernel $P_{Z\mid X}$ induces distributions $P_{Z}=P_{Z\mid X}\circ P_{X}$ and $P_{Z\mid Y=y}=P_{Z\mid X}\circ P_{X\mid Y=y}$. Therefore, we can write
    $\Lambda(Z\to y)$ $=D_{\infty}(P_{Z}\|P_{Z\mid Y=y})$
    $=D_{\infty}(P_{Z\mid X}\circ P_{X}\|P_{Z\mid X}\circ P_{X\mid Y=y})$
@@
  6. (vi)
    Consider the Markov chain XYZX-Y-Z. The kernel $P_{Z\mid Y}$ induces distributions $P_{XZ}=P_{Z\mid Y}\circ P_{XY}$ and $P_{Z}=P_{Z\mid Y}\circ P_{Y}$. Hence, we can write
    $\operatorname*{ess\,sup}_{P_{Z}}\Lambda($ $X\to Z)=D_{\infty}(P_{X}\times P_{Z}\|P_{XZ})$
    $=D_{\infty}(P_{X}\times(P_{Z\mid Y}\circ P_{Y})\|P_{Z\mid Y}\circ P_{XY})$
    $=D_{\infty}(P_{Z\mid Y}\circ(P_{X}\times P_{Y})\|P_{Z\mid Y}\circ P_{XY})$
    $\leq D_{\infty}(P_{X}\times P_{Y}\|P_{XY})$
@@
    $\leq\log\;\operatorname*{ess\,sup}_{P_{X}}\;\frac{f_{Y_{1}}(y_{1})}{f_{Y_{1}\mid X}(y_{1}\mid X)}$ (78)
    $\hskip 50.00008pt+\log\;\operatorname*{ess\,sup}_{P_{X}}\;\frac{f_{Y_{2}\mid Y_{1}}(y_{2}\mid y_{1})}{f_{Y_{2}\mid Y_{1},X}(y_{2}\mid y_{1},X)}$ (79)
    $\leq\log\;\operatorname*{ess\,sup}_{P_{X}}\;\frac{f_{Y_{1}}(y_{1})}{f_{Y_{1}\!\mid\!X}(y_{1}\!\mid\!X)}$ (80)
    $\hskip 50.00008pt+\log\;\operatorname*{ess\,sup}_{P_{X\mid Y_{1}=y_{1}}}\;\frac{f_{Y_{2}\mid Y_{1}}(y_{2}\mid y_{1})}{f_{Y_{2}\mid Y_{1},X}(y_{2}\mid y_{1},X)},$ (81)